Cassini Significant Events 12/05/07 - 12/11/07

December 13, 2007

(Source: Cassini Project)


The most recent spacecraft telemetry was acquired on Tuesday, Dec. 11, from the Madrid tracking complex. The Cassini spacecraft is in an excellent state of health and all subsystems are operating normally. Information on the present position and speed of the Cassini spacecraft may be found on the "Present Position" page at: http://saturn.jpl.nasa.gov/operations/present-position.cfm.

Wednesday, Dec. 5 (DOY 339):

The S39 Science Operations Plan Update (SOPU) official port occurred today. Inputs from participating instrument teams, Spacecraft Operations, and Navigation were merged, and the reports delivered to the teams and AACS for the end-to-end pointing analysis.


On Dec. 3, there was a large spike in Single Bit Errors (SBE) and Double Bit Errors (DBE) on both SSRs. Usually the SBEs run about 25-50 per hour. This spike lasted about 8 hours and got up to about 1850 per hour. DBEs typically run about 2-3 per hour. This spike got up to 6 per hour. SBEs and DBEs are believed to be caused by high energy particle hits. SBEs are autonomously corrected by on-board error detection and correction (EDAC) logic. DBEs in flight software are corrected by autonomous copies of the code from redundant copies in the SSRs. DBEs in data regions of memory are not corrected, but go away with the next writing of data into that area of memory.

Thursday, Dec. 6 (DOY 340):

With Cassini now on the outbound leg of the T38 Titan encounter, a Radio Science (RSS) gravity science experiment, mirroring the one performed as Cassini approached Titan on DOY 338, was executed during the downlink to Earth. For the remainder of the day, Imaging Science (ISS) produced images of Saturn's smaller moons for orbital element determination and observed the ring system, focusing on the diffuse rings and making a mosaic of the ring system as a whole.


The S40 Aftermarket process was completed today. However, Science Planning is still working on reducing Cassini's requested 70m DSN coverage during the Phoenix Entry/Descent/Landing (EDL) contention period. Due to a hiatus in sequence processing over the winter holidays, these changes aren't due until Jan. 9, 2008.

Friday, Dec. 7 (DOY 341):

As Cassini now swings outward beyond Saturn's magnetosphere, the Magnetospheric and Plasma Science (MAPS) instruments resumed their monitoring of solar wind conditions. The Visual and Infrared Mapping Spectrometer (VIMS) observed bright stars to calibrate its sensors, and ISS resumed its monitoring of Saturn's small satellites and produced a ring mosaic.


Orbit Trim Maneuver (OTM) #137 was performed today. This is the cleanup maneuver following the T38 encounter on Dec. 5. The main engine burn began at 11:15 PM PST. Telemetry immediately after the maneuver showed a burn duration of 4.18 seconds, giving a delta-V of 0.68 m/s, as planned. All subsystems reported nominal performance after the maneuver.

Imaging scientists are putting together the story of how the small moons orbiting near the outer rings of Saturn came to be. The moons began as leftover shards from larger bodies that broke apart and filled out their "figures" with the debris that made the rings. For the full story link to:
http://saturn.jpl.nasa.gov/news/press-release-details.cfm?newsID=795

Uplink Operations began sending the Instrument Expanded Block files for S36 today. Seven files for ISS, VIMS, Composite Infrared Spectrometer (CIRS), Optical Navigation, Ultraviolet Imaging Spectrograph (UVIS), and Cassini Plasma Spectrometer (CAPS) were sent. The remaining files will be sent Sunday and the background sequence will go up on Monday. S36 will begin execution on Friday, Dec. 14.


Sunday, Dec. 9 (DOY 343)


Today the SSPS for the SRU-B replacement heater was tripped from OFF to TRIPPED. The SSPS autonomous trip response fault protection algorithm turned it back to OFF several minutes later. There appears to be no correlation between this trip and the SBE and DBE activity seen earlier in the week.

Monday, Dec. 10 (DOY 344)

RSS performed a Saturn Occultation Operational Readiness Test (ORT) today. There will be another ORT performed in the wee hours of the morning on Dec. 16, followed by the occultation itself later that day.

Tuesday, Dec. 11 (DOY 345)

CIRS performed observations for determining the upper troposphere and tropopause temperature of Saturn today.

The DOY-353 Live Movable Block (LMB) kickoff meeting was held today. The meeting addressed the LMB and an additional Live Inertial Vector Propagator update with RSS as the driving instrument. The Science Planning analysis and recommendation is due late tonight with the Go/NoGo meeting to be held tomorrow.